Given two angles that form a linear pair. Prove that thier bisectors are perpendicular.
please give me the statements and reasons. and please keep it as simple as possible. thank you
: Given two angles that form a linear pair. Prove that thier bisectors are perpendicular.
please give me the statements and reasons. and please keep it as simple as possible. thank you

Answer by stanbon(19673)

Given two angles that form a linear pair. Prove that their bisectors are perpendicular.
-----------------
Let the angles be A and B.
A + B = 180 degrees ; def. of linear pair
-----------
Let L1 be a bisector of A forming a1 and a2
Let L2 be a bisector of B forming b1 and b2
Let a1 be ajacent to b1
---------------------
a1=a2 and b1=b2; def. of bisector
a1+a2 = A and b1+b2 = B; def. of bisector
----------------------
2a1 = A and 2b1 = B; substitution
2a1 + 2b1 = 180 ; substitution
---------------
a1 + b1 = 90 ; equals divided by equals are equal
---------------
L1 and L2 are perpendicular ; def of perpendicular
=============================================
Cheers,
Stan H.